Express 2/y2-1 in partial fractions. Hence solve the differential equation 2xdy/dx +1 = y2 given that y= -3 when x= 1, expressing y explicitly in terms of x.


1
Expert's answer
2021-09-20T16:05:47-0400

It follows that "\frac{2}{y^2-1}=\frac{1}{y-1}-\frac{1}{y+1}" is the expression in partial fractions.

Let us solve the differential equation "\frac{2xdy}{dx} +1 = y^2,"  given that "y= -3" when "x= 1." This equation is equvalent to the equation "\frac{2xdy}{dx} = y^2-1," and hencev "\frac{2dy}{y^2-1} = \frac{dx}x." It follows that "\int\frac{2dy}{y^2-1} = \int\frac{dx}x," and hence "\int(\frac{1}{y-1}-\frac{1}{y+1})dy= \ln|x|+\ln|C|." We conclude that "\ln|y-1|-\ln|y+1|=\ln|Cx|," that is "\ln|\frac{y-1}{y+1}|=\ln|Cx|," and hence "\frac{y-1}{y+1}=Cx." Taking into account that "y= -3" when "x= 1," we conclude that "C=\frac{-3-1}{-3+1}=2."

It follows that "y-1=2xy+2x," and hence "y-2xy=2x+1." Consequently, "y=\frac{2x+1}{1-2x}" is the solution of the differential equation.